OpenStudy (anonymous):

Here's another one, please help guys! Solve for x. 2(x - 5)^-1 + 1/x = 0

OpenStudy (dumbcow):

\[\frac{2}{x-5} = -\frac{1}{x}\] \[2x = -(x-5)\] \[2x = -x +5\] \[3x = 5\] \[x = \frac{5}{3}\]

OpenStudy (anonymous):

thanks so much!

OpenStudy (dumbcow):

no problem